OpenStudy (anonymous):

Which of the following is an even function? f(x) = (x – 1)^2 f(x) = 8x f(x) = x^2 – x f(x) = 7

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

D*

OpenStudy (irishboy123):

even function f(x) = f(-x) odd -f(x) = f(-x) test them

OpenStudy (bradely):

Rule for even function is f(-x) = f(x) f(x) = (x – 1)^2 f(-x) = (-x-1)^2 =(x+1)^2 is not equal to f(x) f(x) = 8x f(-x) = -8x is not equal to f(x) f(x) = 7 f(-x) =7 =f(x) f(x) =7 is the even Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

Thank you!